OpenStudy (anonymous):

7/16 = 21/y Directions-Find the value of the letter in each proportion

OpenStudy (anonymous):

7/16 = 21/y y (7/16) = 21 y = (21/7)*16 y = 3*16 y = 48

OpenStudy (anonymous):

7/16=21/y y=(21*16)/7 y=3*16 =48